OpenStudy (anonymous):

what is the answer to negative square root3

OpenStudy (anonymous):

do you mean \[-\sqrt(3)\] ?

OpenStudy (anonymous):

or is the negative inside?

OpenStudy (anonymous):

its like the first one

OpenStudy (anonymous):

I don't know what you mean by answer because sqrt(3) is just a number, if you mean what is the decimal expansion to it then it would be \[-\sqrt(3) = -1.7\]

OpenStudy (anonymous):

-1.73

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

These are called irrational numbers and you can't represent them with finite decimals like let's say 1/2 = 0.5. sqrt(2) = 0. -> a long bunch of decimals. But remember they are just numbers like anything else

OpenStudy (anonymous):

3i

OpenStudy (anonymous):

\[\sqrt{-1}\]=i it an imaginary number.

OpenStudy (anonymous):

\[\sqrt(-3) = i \sqrt(3) \]

OpenStudy (anonymous):

k thanks